PRINTER VERSION Chapter 26, Problem 061 Your answer partially correct. Try again. An objectis 20 am front of advergng lens that has a focal ength o.13 am. How far in front of the lens should the object be placed so that the size as images reduced Dy a factor of 1.8? Click if you would like to Show work for this question Open Show Work ay accessing this Question Assistance, vou will leam while you earn points based on the Point 2 of 10 used Question Attemptse Reserved.
Answer #1

given
object distance d0 = 20cm
focal length f = -13cm
size of the image is reduced by 1.8

we know that
1/do +1/di =1/f
1/20 +1/di = -1/13
di = -7.8 cm
magnification M = -di / do = 7.8/20 = 0.39

(1/1.8)*M = (1/1.8)*0.39
M =0.216
0.216=7.8/do
do =36.11 cm
Answer #2

This will not lead to a correct solution.